OpenStudy (anonymous):

You can use the formula a = h/n is you want to find the batting average a of a batter who has h hits in n times at bat. Solve this equation for h. Then, if Joseph has a batting average of .285 and has been at bat 400 times, how many hits does Joseph have?

OpenStudy (anonymous):

h = a*n

OpenStudy (anonymous):

ok a = .285 n = 400 so .285 = h/400 or 400 x .285 = h =114

OpenStudy (anonymous):

thanks LOVE (: